Calculate the electric field that a test charge will experience at a distance of 2.04 x 10^-3 m from the source charge of 5.02 x 10^-13 C.


1
Expert's answer
2021-10-13T09:07:20-0400

The electric field produced by a point charge is given as follows:


"E = k\dfrac{q}{r^2}"

where "k = 9\times 10^{9}Nm^2/C^2" is the Coulomb constant, "q = 5.02\times 10^{-13}C" is the charge, and "r = 2.04\times 10^{-3}m" is the distance from the source charge. Thus, obtain:


"E = 9\times 10^{9}Nm^2/C^2\cdot \dfrac{5.02\times 10^{-13}C}{(2.04\times 10^{-3}m)^2} \approx 1.09\times 10^3N/C"

Answer. "1.09\times 10^3N/C".